OpenStudy (anonymous):

617.0519 to the nearest tenth

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

wouldn't it be 617.15

OpenStudy (skullpatrol):

$$ 617.0519 =617+\frac{0}{10}+\frac{5}{100}+\frac{1}{1,000}+\frac{9}{10,000}$$

OpenStudy (skullpatrol):

Which is the "tenth" @breannaorona ?

OpenStudy (shamil98):

 you round up since its 5 617.1 is rounded to the nearest tenth